# Grounding LLMs For Robot Task Planning Using Closed-loop State Feedback

### v2

@@ Abstract
Planning algorithms decompose complex problems into intermediate steps that can be sequentially executed by robots to complete tasks. Recent works have employed Large Language Models (LLMs) for task planning, using natural language to generate robot policies in both simulation and real-world environments. LLMs like GPT-4 have shown promising results in generalizing to unseen tasks, but their applicability is limited due to hallucinations caused by insufficient grounding in the robot environment. The robustness of LLMs in task planning can be enhanced with environmental state information and feedback. In this paper, we introduce a novel approach to task planning that utilizes two separate LLMs for high-level planning and low-level control, improving task-related success rates and goal condition recall. Our algorithm, BrainBody-LLM, draws inspiration from the human neural system, emulating its brain-body architecture by dividing planning across two LLMs in a structured, hierarchical manner. BrainBody-LLM implements a closed-loop feedback mechanism, enabling learning from simulator errors to resolve execution errors in complex settings. We demonstrate the successful application of BrainBody-LLM in the VirtualHome simulation environment, achieving a 29% improvement in task-oriented success rates over competitive baselines with the GPT-4 backend. Additionally, we evaluate our algorithm on seven complex tasks using a realistic physics simulator and the Franka Research 3 robotic arm, comparing it with various state-of-the-art LLMs. Our results show advancements in the reasoning capabilities of recent LLMs, which enable them to learn from raw simulator/controller errors to correct plans, making them highly effective in robotic task planning. Demo Video: http://tinyurl.com/2akwhvf2.
@@ 1 Introduction
LLMs, trained on corpora of internet-sourced text, have demonstrated capabilities akin to artificial general intelligence [8]. The inherent world knowledge of LLMs, combined with their in-context learning ability is paving a new direction in robotic task planning. Prior work showed that LLMs can generate step-by-step instructions for complex tasks without any re-training or model parameter updates [17, 19, 43, 24, 52, 3, 36, 42]. Despite promising results in diverse robotic tasks, grounding LLMs in a given environment is still an open problem. Consider the task - “Make me a coffee.” LLMs can decompose this problem into a sequence of steps– ‘1. Walk to fridge,’ ‘2. Grab milk,’ and so on, through to ‘9. Serve cup of coffee.’ Yet, these steps are not entirely executable in a real-world environment with physical constraints. Additional steps like ‘Switch on microwave’ or ‘Open microwave doors,’ are essential for task completion. Moreover, limitations like absence of milk or water should not hinder task execution. Robots must adapt to the environment, refining plans towards successful task completion.
Thus, grounding LLMs in real-world scenes is essential. Incorporating environmental feedback into task planning allows error resolution in real-time, enhancing robot robustness and utility [19]. In this paper, we introduce a novel planning algorithm that aims at mitigating two issues in existing methods: i) Our approach uses a simple prompting framework, with clear distinction of planning, feedback and execution components to avoid using expert defined heuristics and carefully constructed world models and ii) Our method eliminates human intervention through utilizing raw error messages from simulators and controllers to guide an LLM planner for robust task execution, enhancing autonomy. Our contributions are:
@@
    A novel planning algorithm that uses a Two-LLM system (Brain-LLM and Body-LLM) to derive executable actions from natural language instructions, leveraging a closed-loop state feedback mechanism for error resolution (Figure 1).
@@
    Improving task-oriented success rate by 29% average over existing state-of-the-art techniques in the VirtualHome Embodied Control environment (using a dataset of 80 tasks). BrainBody-LLM on average completes 72% of all goal conditions for a given task, improving over existing methods.
@@ 2 Background and Related Work
2.1 Foundational Models
Designing deep learning systems capable of comprehending and generating natural language has long been an attractive yet challenging task in the AI community. The emergence of transformers [47] and the success of pre-trained language models such as BERT [11] have paved the way for more advanced foundational models known as Large Language Models [45, 4, 1]. These LLMs are trained on extensive textual data using masked language modeling and autoregressive text prediction. The discovery of in-context learning capabilities, a feature absent in earlier models, has laid the foundation for various applications of LLMs [9]. In this study, we exploit these capabilities, particularly prompt engineering, which requires no gradient updates in the foundational model.
2.2 LLMs in Robotic Task Planning
@@
The real world knowledge contained within LLMs can be utilized in robots for common sense reasoning and generating language conditioned policies for task execution. Earlier works trained smaller LLMs like GPT-2 for agent task planning, treating it as a translation problem from natural language instructions to high level plans [20, 27]. More recently, training LLMs in a multi-task setup for embodied control has demonstrated robust performance in planning, video captioning, video QnA and multi-turn dialogue [29]. Extremely large Vision Language Models (VLMs) trained on diverse robot demonstration data can be used to directly convert vision-language input modalities to physical actions in the real world [6, 7, 2, 13]. Training VLMs is highly resource intensive, and accurate real world deployment requires customization through further environment specific finetuning [49].
@@
Foundational models after GPT-3 have demonstrated excellent in-context learning abilities, which enables learning from examples in natural language without the need for expensive training and model updates. By choosing a suitable language prompt and some domain examples, LLMs can break down complex tasks into executable steps complying agent-environment constraints for high level planning [17]. LLMs can be used to learn from structured data such as Planning Domain Definition Language (PDDL) or Hierarchical Linear Temporal Logic (HLTL) and combined with heuristic search based planners [39, 26]. LLM generated PDDL world models can be used with domain independent planning models in robotic task planning [16, 51]. Constructing PDDL world models is difficult in real world scenarios, where object and environment states are often revealed in runtime thus requiring human assistance for effective policy exploration.
LLM generated plans sometimes consist of hallucinatory output, when the planner uses objects or actions that do not exist in the environment. This problem is mitigated by downstream filtering mechanisms that improve the correctness of LLM generated plans. [33] uses a conformal prediction scheme to rank LLM generated plans for a current task. High level plans can be combined with motion planning and RL control policies for physical execution [10]. Visual perception greatly aids LLM generated sequence of skills through geometric conditioned policy selection [22]. Combining LLMs with visual perception through VLMs to construct 3D value maps can be used by high level planners for effective grounding in real world environments [18]. Subsequent research has demonstrated application of such systems in navigation tasks through visual grounding [40, 53] or incrementally generated 3D scene graphs [34, 35].
@@
Decision making skills of goal driven LLM agents can be improved with runtime feedback [37] or affordance function based grounding [3]. LLMs deployed in robots can incorporate feedback through natural language, and can use it to correct erroneous plans to enhance overall task success rate [24, 46]. Feedback from human intervention or simulator messages can improve real world alignment [16]. Methods that use environmental feedback for task planning can be classified into two broad categories: static and dynamic planners. Static planners use feedback to check whether the robot’s environment satisfies the necessary conditions outlined in the generated plan, thus preventing execution errors. In this case, the feedback does not alter the generated plans but improves task oriented success rate [24, 36]. On the other hand, dynamic planners use feedback from the system to assert necessary conditions and to alter the plans, improving their execution [19, 43, 52, 42]. ProgPrompt, a well-known technique, was one of the pioneers of LLM-based planning for robotic tasks [36]. Their algorithm used LLMs to break down complex high-level instructions into sequential steps by writing Python functions utilizing a known set of API skills. Assert statements (if-else) checked the existence of certain preconditions (e.g., whether the microwave door is open or the cup is on the table). This static planning mechanism improved both the success rate and task execution, as feedback from assert statements was utilized by LLMs to create robust plans. Our approach distinguishes itself from ProgPrompt and other conventional dynamic planners by employing a Two-LLM system designed to assimilate error messages from a simulator and environmental states in tandem for enhanced task planning. This allows our planner to understand errors in natural language and generate corrected plans for successful execution. Thus, we move one step further from ProgPrompt by replacing simple assertion-based precondition checks with natural language and real-time feedback. To the best of our knowledge, we are the first to implement a dynamic planning algorithm for the VirtualHome Simulator and Franka Research 3 arm. More differences with other LLM based planning algorithms are described in Section 3.4.
LLM generated plans are often in natural language and need to be converted to suitable action primitives for parsing and execution [51]. These action primitives must follow strict syntactic rules, as errors in these control statements can lead to downstream execution problems. A popular solution involves designing control functions in known programming languages like Python, which are then translated into rule-based action primitives [24, 36]. Our method uses simple action statements as execution primitives, but can be conditioned to output python code for robot execution, similar to previous studies in this area.
@@
2:$T$triangleright Task description
3:$K$triangleright Max feedback loops
@@
5:$\phi(\text{task},\text{feedback})$triangleright Brain-LLM
6:$\theta(\text{high level plan step})$triangleright Body-LLM
7:$\pi(\text{low level action})$triangleright Actuator
8:$\rho(\text{action result})$triangleright Simulator/Perception
@@
10:$k\leftarrow 0$triangleright feedback loop counter
11:$f\leftarrow$none $\triangleright$ no feedback in the beginning
12:$hlp\leftarrow\phi(T,f)$triangleright high Level Plan
13:$i\leftarrow 0$
@@
15:    $step\leftarrow hlp[i]$triangleright ith step of high level plan
16:    $\delta\leftarrow\theta(step)$triangleright low level Action
@@
21:                 $f\leftarrow\rho(action\_result)$triangleright feedback
@@
25:                 $i\leftarrow i+1$triangleright go to next step
@@
27:         $\triangleright$ action is successful
28:             $i\leftarrow i+1$triangleright go to next step
@@
30:    $\triangleright$ no primitive for the current action
31:         $i\leftarrow i+1$triangleright go to next step
@@ 3 Our Proposal: BrainBody-LLM
The human brain is a sophisticated information processing system comprising a network of billions of neurons. Artificial neural networks aim to mimic brain functions, but a significant difference lies in the information processing method: the human brain functions continuously, in contrast to the discrete operation of computer algorithms [5, 21]. To emulate critical cognitive functions of the human brain, particularly in terms of feedback mechanisms and adaptability, we developed the BrainBody-LLM algorithm.
@@
Our algorithm consists of two LLMs, each with a distinct contribution to the overall task execution pipeline. The Brain-LLM is designed to decompose a given task into high-level execution plans in natural language. High-level plans are converted by the Body-LLM into low level robot control commands, with a pre-defined syntax for environmental execution. During runtime, any error messages provided by the simulation environment, motion planning controller or human feedback can be relayed to the Brain-LLM, which then generates an updated plan from the current step to complete the original task. The updated plan is subsequently converted to control statements by the Body-LLM. Our method uses iterative planning and feedback to create a closed loop pipeline. Our complete framework of integrating two LLMs is described in Algorithm 1. All LLM prompts in our approach include in-context learning examples that inform the planner about the environmental constraints (e.g., the microwave is closed or the TV is on) and provide accurate, manually curated task-plan example pairs. BrainBody-LLM uses these examples to uncover patterns and understand object-action relationships to generate plans for unseen tasks. Moreover, feedback through error messages further grounds the planner in the real world by informing it about erroneous steps and explaining the reasons for the failure of particular steps.

### v3

@@ Abstract
[…]
@@ 1 Introduction
LLMs, trained on corpora of internet-sourced text, have demonstrated capabilities akin to artificial general intelligence [bubeck2023sparks]. The inherent world knowledge of LLMs, combined with their in-context learning ability is paving a new direction in robotic task planning. Prior work showed that LLMs can generate step-by-step instructions for complex tasks without any re-training or model parameter updates [huang2022language, pmlr-v205-huang23c, Sun2023AdaPlannerAP, 10160591, yao2023react, Ahn2022DoAI, progprompt, song2023llm, Kambhampati2024LLMsCP]. Despite promising results in diverse robotic tasks, grounding LLMs in a given environment is still an open problem. Consider the task - “Make me a coffee.” LLMs can decompose this problem into a sequence of steps– ‘1. Walk to fridge,’ ‘2. Grab milk,’ and so on, through to ‘9. Serve cup of coffee.’ Yet, these steps are not entirely executable in a real-world environment with physical constraints. Additional steps like ‘Switch on microwave’ or ‘Open microwave doors,’ are essential for task completion. Moreover, limitations like absence of milk or water should not hinder task execution. Robots must adapt to the environment, refining plans towards successful task completion.
Thus, grounding LLMs in real-world scenes is essential. Incorporating environmental feedback into task planning allows error resolution in real-time, enhancing robot robustness and utility [pmlr-v205-huang23c]. In this paper, we introduce a novel planning algorithm that aims at mitigating two issues in existing methods: i) Our approach uses a simple prompting framework, with clear distinction of planning, feedback and execution components to avoid using expert defined heuristics and ii) Our method can integrate feedback from simulators, controllers or human intervention to guide an LLM planner for robust task execution, enhancing autonomy. Our contributions are:
@@
    A novel planning algorithm that uses a Two-LLM Agentic framework (Brain-LLM and Body-LLM) to derive executable actions from natural language instructions, leveraging a closed-loop state feedback mechanism for error resolution (Figure 1).
@@
    Improving task-oriented success rate by 15% average over existing state-of-the-art techniques in the VirtualHome Embodied Control environment (using a dataset of 80 tasks). BrainBody-LLM on average completes 81% of all goal conditions for a given task.
@@ 2 Background and Related Work
2.1 Training LLMs for Robotic Policy Generation
@@
The real world knowledge contained within LLMs can be utilized in robots for common sense reasoning and generating language conditioned policies for task execution. Earlier works trained smaller LLMs like GPT-2 for agent task planning, treating it as a translation problem from natural language instructions to high level plans [jansen-2020-visually, micheli-fleuret-2021-language]. More recently, training LLMs in a multi-task setup for embodied control has demonstrated robust performance in planning, video captioning, video QnA and multi-turn dialogue [mu2024embodiedgpt]. Extremely large Vision Language Models (VLMs) trained on diverse robot demonstration data (comprising frame aligned RGB images and action vectors) can be used to directly convert vision-language input modalities to physical actions in the real world [rt12022arxiv, rt22023arxiv, Ahn2022DoAI, driess2023palme, kim2023lalm, belkhale2024rt]. Training VLMs is highly resource intensive, and accurate real world deployment requires customization through further environment specific training [Wake2023GPT4VisionFR]. Although methods like OpenVLA [kim24openvla] facilitate LoRA-based fine-tuning to reduce computational requirements, they cannot be deployed in zero-shot settings for novel environments without first collecting task-specific demonstrations and undergoing re-training.
2.2 Prompting LLMs for Task Planning
@@
Foundational models after GPT-3 have demonstrated excellent in-context learning abilities, which enables learning from examples in natural language without the need for expensive training and model updates. By choosing a suitable language prompt and some domain examples, LLMs can break down complex tasks into executable steps complying agent-environment constraints for high level planning [huang2022language, wang2023voyager, mahadevan_gen]. [yao2023react] introduces an LLM framework for decision making in simulation environments. LLMs can be used to learn from structured data such as Planning Domain Definition Language (PDDL) or Hierarchical Linear Temporal Logic (HLTL) and combined with heuristic search based planners [silver2022pddl, luo2023obtaining]. LLM generated PDDL world models can be used with domain independent planning models in robotic task planning [guan2023leveraging, Xie2023TranslatingNL]. Constructing PDDL world models is difficult in real world scenarios, where object and environment states are often revealed in runtime thus requiring human assistance for effective policy exploration. LLM generated plans are often in natural language and need to be converted to suitable action primitives for parsing and execution [Xie2023TranslatingNL]. These action primitives must follow strict syntactic rules, as errors in these control statements can lead to downstream execution problems. A popular solution involves designing control functions in known programming languages like Python, which are then translated into rule-based action primitives [10160591, progprompt]. Our method uses simple action statements as execution primitives, but can be conditioned to output python code for robot execution, similar to previous studies in this area.
2.3 Grounding LLMs to prevent Hallucination
@@
LLM generated plans sometimes consist of hallucinatory output, when the planner uses objects or actions that do not exist in the environment. This problem is mitigated by downstream filtering mechanisms that improve the correctness of LLM generated plans. [knowno2023] uses a conformal prediction scheme to rank LLM generated plans for a current task. High level plans can be combined with motion planning, active preference learning and RL control policies for physical execution [wang2024apricot, dalal2024psl]. Visual perception greatly aids LLM generated sequence of skills through geometrically conditioned policy selection [Lin2023, shah2024bumble]. Combining LLMs with visual perception through VLMs to construct 3D value maps can be used by high level planners for effective grounding in real world environments [huang2023voxposer]. Subsequent research has demonstrated application of such systems in navigation tasks through visual grounding [shah2022lmnav, zhou2023navgpt] or incrementally generated 3D scene graphs [rana2023sayplan, Rajvanshi_Sikka_Lin_Lee_Chiu_Velasquez_2024]. Decision making skills of goal driven LLM agents can be improved with runtime feedback [shinn2023reflexion] or affordance function based grounding [Ahn2022DoAI]. LLMs deployed in robots can incorporate feedback through natural language, and can use it to correct erroneous plans to enhance overall task success rate [10160591, Valmeekam2023OnTP]. Feedback from human intervention or simulator messages can improve real world alignment [guan2023leveraging]. Methods that use environmental feedback for task planning can be classified into two broad categories: static and dynamic planners. Static planners use feedback to check whether the robot’s environment satisfies the necessary conditions outlined in the generated plan, thus preventing execution errors. In this case, the feedback does not alter the generated plans but improves task oriented success rate [10160591, progprompt]. On the other hand, dynamic planners use feedback from the system to assert necessary conditions and to alter the plans, improving their execution [pmlr-v205-huang23c, Sun2023AdaPlannerAP, yao2023react, song2023llm]. ProgPrompt, a well-known technique, was one of the pioneers of LLM-based planning for robotic tasks [progprompt]. Their algorithm used LLMs to break down complex high-level instructions into sequential steps by writing Python functions utilizing a known set of API skills. Assert statements (if-else) checked the existence of certain preconditions (e.g., whether the microwave door is open or the cup is on the table). This static planning mechanism improved both the success rate and task execution, as feedback from assert statements was utilized by LLMs to create robust plans. Our approach distinguishes itself from ProgPrompt and other conventional dynamic planners by employing a Two-LLM system designed to assimilate error messages from a simulator and environmental states in tandem for enhanced task planning. This allows our planner to understand errors in natural language and generate corrected plans for successful execution. Thus, we move one step further from ProgPrompt by replacing simple assertion-based precondition checks with natural language and real-time feedback. To the best of our knowledge, this work is the first to implement a dynamic planning algorithm for both the VirtualHome Simulator and the Franka Research 3 robotic arm.
2.4 Closed loop LLM Planning
Our work derives inspiration from previous methods in utilizing feedback by back-prompting LLMs for improved robotic task planning [Valmeekam2023OnTP, wang2023describe, skreta2023errors]. A hierarchical planner, which breaks down complex task instructions to intermediate planning and action primitives has been explored before [song2023llm]. However, we design our system to ensure minimal domain customization, and demonstrate applicability in both simulation and real world setups. rana2023sayplan uses two distinct LLMs for 3D scene graph search and iterative re-planning using simulator feedback. Our approach does not rely on scene graphs, with planning and error resolution performed by the Brain-LLM and action mapping performed by Body-LLM. pmlr-v205-huang23c implements a closed loop feedback integrating visual perception and language conditioned robotic skills. In their approach, the LLM planner has access to vision models providing information about the environment, and it further disambiguate planning confusion by asking questions to a human operator. Our contributions lie in an effective and autonomous planning approach that eliminates any human-in-the-loop scenario by reinforcing feedback through error messages. There are a number of works that integrate visual grounding in planning with VLMs for object manipulation [Wu2023TidyBotPR, Liu2024OKRobotWR, huang2023voxposer, chen2023llm, liu2023reflect] and navigation tasks [shah2022lmnav, zhou2023navgpt], and we leave such expansion of our planning methodology for future work. Our work can also be integrated with skill learning techniques through human-guided imitation learning as demonstrated in [parakh2023humanassisted].
A two-LLM system for generating symbolic and geometric relationships for task and motion planning has shown effectiveness in robotic environments [10342169]. Our method distinguishes itself by avoiding any human expert intervention or customization for filtering, allowing environmental feedback and error messages to automatically guide the planning LLM towards an accurate and correct real world action. Traditional simulators—human-crafted world models—have long supplied the environmental feedback and state information used for robot planning. Recent advances in camera-based perception, however, offer far richer visual context for LLM-driven planners [molmo2024, ravi2025sam]. Since our algorithm is perception-agnostic, it incorporates these improvements without modification, providing a straightforward yet highly scalable framework for end-to-end, language-guided robot control. Silver_Dan_Srinivas_Tenenbaum_Kaelbling_Katz_2024 uses two LLMs for planning and correction using execution feedback, but relies on PDDL models for effective grounding, whereas our approach eliminates this dependency and can be used in real world experiments where objects and environment states are not known during task planning. The closest to our work is [li2023interactive], which also uses an LLM for high level step by step task planning, followed by another LLM for converting plans to low level functional APIs for robot control. Their method is capable of re-planning based on human-in-the-loop feedback, while ours uses task errors (which can be through simulator or a human feedback) to re-plan.
@@ 2.4 Closed loop LLM Planning
2:$T$ $\triangleright$ Task description
3:$K$ $\triangleright$ Max feedback loops
@@
[…]
@@
[…]
@@
15:  $step\leftarrow hlp[i]$ $\triangleright$ ith step of high level plan
16:  $\delta\leftarrow\theta(step)$ $\triangleright$ low level Action
@@
21:      $f\leftarrow\rho(action\_result)$ $\triangleright$ feedback
@@
25:      $i\leftarrow i+1$ $\triangleright$ go to next step
@@
[…]
28:     $i\leftarrow i+1$ $\triangleright$ go to next step
@@
[…]
31:    $i\leftarrow i+1$ $\triangleright$ go to next step
@@ 3 Our Proposal: BrainBody-LLM
The human brain is a sophisticated information processing system comprising a network of billions of neurons. Artificial neural networks aim to mimic brain functions, but a significant difference lies in the information processing method: the human brain functions continuously, in contrast to the discrete operation of computer algorithms [korteling_et_al, aimone_et_al]. To emulate critical cognitive functions of the human brain, particularly in terms of feedback mechanisms and adaptability, we developed the BrainBody-LLM algorithm.
@@
Our algorithm consists of two LLMs, each with a distinct contribution to the overall task execution pipeline. The Brain-LLM is designed to decompose a given task into high-level execution plans in natural language. High-level plans are converted by the Body-LLM into low level robot control commands, with a pre-defined syntax for environmental execution. During runtime, any error messages provided by the simulation environment, motion planning controller or human feedback can be relayed to the Brain-LLM, which then generates an updated plan from the current step to complete the original task. The updated plan is subsequently converted to control statements by the Body-LLM. Our method uses iterative planning and feedback to create a closed loop pipeline. Our complete framework of integrating two LLMs is described in Algorithm 1. The framework begins by passing the task description $T$ and initial empty feedback $f$ to the Brain-LLM $\phi$, producing a high-level plan hlphlp. The loop index $i$ iterates through each step of this plan, and the Body-LLM $\theta$ converts the current step into a low-level action $\delta$, which the actuator policy $\pi$ executes to yield an action_resultaction\_result. If the result is not a success, the perception/simulator module $\rho$ extracts new feedback $f$; provided the feedback-loop counter $k$ is below the maximum $K$, this feedback is sent back to $\phi$ for on-the-fly re-planning, otherwise the algorithm proceeds to the next step. The procedure continues until every plan step succeeds or no further primitives are available, giving a simple yet flexible two-LLM framework that couples high-level reasoning with low-level control while supporting up to $K$ adaptive replanning cycles. All LLM prompts in our approach include in-context learning examples that inform the planner about the environmental constraints (e.g., the microwave is closed or the TV is on) and provide accurate, manually curated task-plan example pairs. BrainBody-LLM uses these examples to uncover patterns and understand object-action relationships to generate plans for unseen tasks. Moreover, feedback through error messages further grounds the planner in the real world by informing it about erroneous steps while explaining reasons for the failure.
State Representation: Our algorithm assumes access to a continuous state representation of the environment during task execution. State information is utilized to recover from execution failures, enabling the robot to realign and complete the task. We represent the scene at time $t$ by the state set
$S_{t}=\bigl(\mathcal{O}_{t},\,\mathcal{A}_{t},\,\mathcal{R}_{t}\bigr),$
where $\mathcal{O}_{t}=\{o_{1},\dots,o_{|\mathcal{O}_{t}|}\}$ denotes the objects currently present, $\mathcal{A}_{t}=\{a_{1},\dots,a_{|\mathcal{A}_{t}|}\}$ the robot actions executable in this state, and $\mathcal{R}_{t}=\{(s{:}p{:}o)_{k}\}_{k=1}^{|\mathcal{R}_{t}|}$ the relational triplets describing object properties and pairwise relations—for example, $(\textit{fridge}:{\textit{isOpen}}:{\textit{true}})$ or $(\textit{cup}:{\textit{on}}:{\textit{table}})$. In simulation environments, such states are typically available through built-in perception modules. In real-world trials, these states can be constructed using visual perception techniques. In this work, the terms ‘environmental information’ and ‘state representation’ are used interchangeably. For instance, [guan2024task] propose leveraging vision-language models (VLMs) to identify undesirable robot behaviors from demonstrations, providing feedback to refine planning. However, accumulating such states in real-time can create bottlenecks in the deployment of large language model (LLM) algorithms. Recent advancements in real-time visual tracking algorithms, such as those by [yang2024samurai], offer potential solutions to mitigate these issues. Our contribution lies in developing an algorithm utilizing these states to enhance planning, with the accumulation of state information being outside the scope of this work.